Session: Oct 2007 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Hartebeesthoek (Jonathan Quick at Tue Nov 6 01:06:46 2007) Success :
No known problems

Torun (Kaz Borkowski at Tue Nov 6 06:23:45 2007) Success :
No known problems.

Onsala (Michael Lindqvist at Tue Nov 6 06:31:11 2007) Failure : WIND
Could not observe due to high wind.

Medicina (Andrea Orlati at Tue Nov 6 09:10:30 2007) Success :
No known problems

Jodrell_Bank (Alastair Gunn at Tue Nov 6 10:05:03 2007) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Tue Nov 6 10:06:01 2007) Minor failures : ANT POIN
Telescope control software kept sending random offset values, first noted at 21:36 UT (reason unknown). Restarted FS with new schedule including corrected offsets at 23:39 UT. There may be some sporadic loss of fringes due to incorrect pointing offsets. No other known problems.

Effelsberg (A. Polatidis at Tue Nov 6 11:15:02 2007) Success :
Lost scan 0538-0538 UT. Bad scan_check at the last two scans, otherwise no problems.

Noto (P. Cassaro at Tue Nov 6 15:08:04 2007) Minor failures :
The phase-cal was on. I have not received the PI communication in time.